1.How many mL of 0.250 M barium nitrate are required to precipitate as barium sulfate all the sulfate ions from 50 mL of 0.350 M aluminum sulfate?

3 Ba(NO3)2 (aq)           +       Al2(SO4)3 (aq)     -->   3 BaSO4 (s)  + 2 Al(NO3)3 (aq)


  1. 2.5 mL of nitric acid, HNO3, are required to react with 50 mL of a 1.25 M Barium hydroxide, Ba(OH)2 solution:

2 HNO3 (aq)           +      Ba(OH)2 (aq) -->    2 H2(aq)    +      Ba(NO3)2 (aq)

Determine the concentration (Molarity) of the nitric acid solution (HNO3).


1
Expert's answer
2022-02-11T11:32:01-0500

Molar mass of barium nitrate 261.337

0.25×261.337=65.33g

Molar mass of aluminum sulfate = 342.15 g/mol

0.35×342.15=119.8

= 119.8/65.33=1.83moles